A satellite is in an elliptical orbit at altitudes ranging from 220 km to 900 km. At its highest point, its moving at 7.43 km/s. How fast is it moving at the lowest point?
Answer #1

let

r1 = Re + h1

= 6.37*10^6 + 220*10^3

= 6.59*10^6 m

r2 = Re + h2

= 6.37*10^6 + 900*10^3

= 7.27*10^6 m

v2 = 7.43 km/s

v1 = ?

let m is the mass of satellite.

apply conservation of angular momentum

m*v1*r1 = m*v2*r2

==> v1 = v2*(r2/r1)

= 7.43*(7.27*10^6)/(6.59*10^6)

= 8.20 km/s <<<<<<<<<----------------Answer